OpenStudy (anonymous):

If dy/dx is 2x-7 in an initial value problem, and y(2) = 0, how do I go about solving this? I tried plugging in 0 for x, but I don't think it's quite that simple...

OpenStudy (anonymous):

u need to integrate 2x-7. do u know how to do that?

OpenStudy (anonymous):

is that finding the antiderivative?

OpenStudy (anonymous):

yes :d

OpenStudy (anonymous):

ok,so it's x^2 - 7x. Um, Idk what to do after. Plug 2 in?

OpenStudy (radar):

What are you solving for?

OpenStudy (anonymous):

It just says "solve the initial value problem" and the answer is an equation, but I have no idea how it got there. Answer is: x^2 -7x +10. So I guess I'm looking for C using y(2)=0?

OpenStudy (anonymous):

from x^2 - 7x +C from integrating original prob.

OpenStudy (radar):

So you are to solve for C Hmmm. 0=2^2 -14 + C substituting 2 for x 0=4-14+c C=10 Would that work?

OpenStudy (anonymous):

Oh wow, that works! Thanks! :)

OpenStudy (radar):

I hope that is the correct procedure, and you are certainly welcome. good luck with your studies.